OpenStudy (anonymous):

lim x->0 (sin2x/x)=?

OpenStudy (anonymous):

@Loser66

OpenStudy (campbell_st):

you need 2 x in the denominator . That way you can use the sin(u)/u limit so increase the denominator to 2x... and multiply by 2 as 1/2 * 2 = 1 \[ 2 \times \lim_{x \rightarrow 0} \frac{\sin(2x)}{2x} \]